OpenStudy (anonymous):

Write the equation in rectangular form: 3=rcos(theta-pi/3) I need help figuring this out

OpenStudy (anonymous):

Your basic conversion equations are \[\begin{cases} x=r\cos\theta\\ y=r\sin\theta\\ x^2+y^2=r^2\\ \tan\theta=\dfrac{y}{x}\end{cases}\] The first two give rise to the other two. As a first step, I'd recommend writing \(\cos\left(\theta-\dfrac{\pi}{3}\right)\) in terms of \(\cos\theta\). Do you recall a useful identity you can use to accomplish this?

OpenStudy (anonymous):

i onlt know of the\[\cos(\theta-\left(\begin{matrix}2 \\ \pi\end{matrix}\right)\]

OpenStudy (anonymous):

Uh, not sure what you're referring to. Your notation doesn't make sense to me. I'm referring to the following identity: \[\cos(x\pm y)=\cos x\cos y\mp\sin x\sin y\] In this case, you have \[\begin{align*}\cos\left(\theta-\dfrac{\pi}{3}\right)&=\cos \theta\cos \frac{\pi}{3}+\sin x\sin \frac{\pi}{3}\\ &=\frac{1}{2}\cos\theta+\frac{\sqrt3}{2}\sin\theta \end{align*}\]

OpenStudy (anonymous):

do you know exactly what that identity is called? because i had no idea that existed,our teacher made no reference to it that i know of

OpenStudy (anonymous):

I think the name is just angle sum/difference identity.

OpenStudy (anonymous):

aah ok, well i will look into it later, i doubt that this type of question will be on my test, i was just trying to do it for tries I gtg thanks alot, and is that how the final answer looks?

OpenStudy (anonymous):

You're welcome, but no, you still have yet to convert to rectangular. I only suggested this so that you can more easily manage the conversion.

OpenStudy (anonymous):

The only thing you're missing is substituting the obtained expression into the original equation. \[3=r\cos\left(\theta-\frac{\pi}{3}\right)~~\Rightarrow~~3=\frac{1}{2}r\cos\theta+\frac{\sqrt3}{2}r\sin\theta\] Now you just use the conversions I listed earlier, and you get \[3=\frac{1}{2}x+\frac{\sqrt3}{2}y\] This would be the answer. Notice there are no \(r\)'s or \(\theta\)'s anywhere in the final answer.

OpenStudy (anonymous):

oh ok, cool that does make alot more sense. It turns out i did know that identity i had just completely forgotten it existed. Thanks alot